OpenStudy (anonymous):

y varies directly as the square of x. When x = 4, y = 8. Find x when y = 32. x = ±4 x = ±6 x = ±8 x = ±12

OpenStudy (lgbasallote):

y varies directly as the square of x so y = kx^2 when x = 4 then y = 8 y = kx^2 8 = k(4)^2 8 = 16k 8/16 = k 1/2 = k what is x when y = 32 y = kx^2 32 = (1/2)x^2 (32 x 2) = x^2 \(\small \sqrt{32 \times 2}\) = x does that help?